# Thunderdome: Timelock-Free Rationally-Secure Virtual Channels

### v3

@@ 1.2 Our contribution
  • We formalize the security properties of Thunderdome and conduct a security analysis within the honest/Byzantine model (Section 6). Specifically, we demonstrate that our protocol is secure under the assumption of distrusting participants, where $f$ out of $13f+$ 1 wardens in each payment channel are Byzantine, and the remainder are honest.
@@ 2.2 Brick channel
Channel structure. In addition to the two primary parties, Alice and Bob, a Brick channel includes a committee of $13f+$ 1 wardens, where up to $f$ wardens can behave maliciously (Byzantine). The basic structure of the Brick channel is illustrated in Fig. 1.
@@
Brick open. Once Alice and Bob agree to open a Brick channel, they broadcast the channel information to the wardens. Each party can consider the channel open only after receiving at least $12f+$ 1 signed acknowledgments from the wardens – a quorum certificate. The threshold of $12f+$ 1 ensures safety in asynchronous communication networks.
Brick update. Each state generated by Alice and Bob is hashed and signed by both parties. The state is then assigned a sequence number to indicate its order. This sequence number, along with the hashed state, is signed and broadcast to the wardens. Once at least $12f+$ 1 wardens have signed the sequence number, the state is considered valid, and the parties can execute it.
Brick close. The Brick channel closing protocol has two scenarios: optimistic and pessimistic. In the optimistic case, both parties are online and responsive, allowing them to collaboratively generate and sign a closing request, which is then published on-chain. In the pessimistic case, one party, say Alice, may be offline for an extended period and unresponsive to Bob’s close request. To prevent the channel from being indefinitely locked, the Brick protocol allows Bob to close the channel unilaterally. However, Bob alone does not have a valid closing request, as during the channel update, the parties only exchange signatures on the hashed state, which is insufficient to close the channel. To close unilaterally, Bob initiates a closing request on-chain, prompting the wardens to publish the latest stored sequence number. The valid closing state is defined by the highest sequence number signed by Alice, Bob, and at least one warden. To incentivize honest behavior from the wardens, Brick employs a punishment mechanism based on proofs-of-fraud. A proof-of-fraud consists of a warden’s signature on an update with a higher sequence number than the one they submitted for the closing request. Bob collects these signatures during the channel update process. If a valid proof-of-fraud is provided, Bob can claim the corresponding warden’s collateral on-chain. According to Brick’s security analysis, if the aggregate collateral of wardens submitting outdated states exceeds the channel balance, Bob has no incentive to close the channel incorrectly. Instead, he would claim the wardens’ collateral and award the entire channel balance to his counterparty. Specifically, for a Brick channel with $v$ coins as the balance, each of the $13f+$ 1 wardens must deposit at least $\frac{v}{f}$ coins on-chain to ensure security.
@@ 3.2 Threat models
Byzantine security model. In this model, we assume each committee consists of $13f+$ 1 wardens, of which $f$ may be Byzantine, while the rest are honest and follow the protocol as specified. We note that the protocol’s security can be extended to scenarios where the number of wardens differs as long as the Byzantine ratio remains no larger than $\frac{f}{3f+1}$ (i.e., the total number of wardens may vary depending on the setup, but the Byzantine proportion must be maintained). In addition to wardens, the security model also accounts for cases where the main Thunderdome parties may be Byzantine, ensuring security for any honest participant. This is the weakest meaningful assumption, as collusion between all parties against a single honest party could occur; however, considering all parties deviating from the protocol holds no value since security properties apply only to honest participants. Byzantine nodes can arbitrarily deviate by delaying, crashing, or signing/publishing incorrect states or messages. However, they cannot drop honest messages or forge signatures, as these actions violate network and cryptographic assumptions.
@@ 4 Protocol Overview
Naive design. In a naive design of Thunderdome, we follow the same rationale as regular TPC protocols like Brick. When opening the virtual channel, the wardens of each TPC lock a portion of their collateral into Thunderdome. For example, if 2 out of 10 coins in a TPC are locked in the virtual channel, each warden locks 20% of their collateral during the funding transaction. The update phase of Thunderdome operates similarly to Brick. However, the closing phase differs during collaborative closing: Thunderdome cannot be closed solely by Alice and Bob; Ingrid’s agreement is also required. For unilateral closing, the smart contract only requires a quorum certificate ($12f+$ 1 signatures) from the closing party’s warden committee (e.g., $C_{A}$ or $C_{B}$).
@@
Once we establish how smart contracts should handle states published by wardens, the next question is how many on-chain publications are necessary, given that asynchrony only guarantees eventual message delivery. If an updated state is confirmed with only $12f+$ 1 signatures from each warden committee, there could still be up to $f$ (honest) wardens who have not received the latest state, even after the protocol was closed (incorrectly). To mitigate this, we require the smart contract to wait for at least $f+1$ publishments from the corresponding committee before closing the channel. This ensures that at least one warden with locked collateral knows the latest state.
@@
Thus, we require at least $12f+$ 1 wardens from each committee, $C_{A}$ and $C_{B}$, to publish their closing state in Thunderdome’s unilateral closing protocol. Combined with the quorum certificates required to update the Thunderdome state, we ensure that at most $f$ wardens per committee can be slow. By waiting for $12f+$ 1 publications from accountable wardens, at least $f+1$ publications can be punished if they submit outdated information. Given that each warden locks $v/f$ coins in Thunderdome, the total collateral that can be claimed per channel in case of fraud is $(f+1)\times\frac{v}{f}>v$.
@@ 5.1 Thunderdome Open
0:  Parties $A$, $I$ and $B$, TPC committees $C_{A}$ and $C_{B}$, initial virtual channel state $s_{1}$, total channel balance $v$ and contract information $ci$.
@@
1:  $A$, $I$ and $B$ combine wardens from $C_{A}$ and $C_{B}$ to form virtual channel committee $C_{V}=C_{A}\cup C_{B}$. /*Parties reach agreement on the open state*/
2:  $A$ and $B$ generate two pre-register transaction with their signatures $TX_{r}^{\prime}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A}/\sigma_{B}\}$ and send to $I$.
3:  After receiving two transactions, $I$ responds with $TX_{r}^{\prime\prime}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A}/\sigma_{B},\sigma_{I}\}$.
4:  $A$ and $B$ exchange responses of $I$ and generate final register transaction with all partis’ signatures $TX_{r}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A},\sigma_{B},\sigma_{I}\}$.
@@
6:  $A$ and $I$ ($I$ and $B$) update payment channels according to initial virtual channel state $s_{1}$.
@@
The final step in opening Thunderdome, after all parties reach an agreement, is to "virtually lock" the coins for the virtual channel. Unlike payment channels, which require publishing funding transactions on-chain, the parties of both underlying payment channels—Alice and Ingrid, and Ingrid and Bob—update their respective payment channels to lock coins for use in the virtual channel. For example, if the payment channel state between Alice and Ingrid is $(a,b)$, and Alice wants to open a virtual channel with Bob with an initial state of $(c,d)$, then Alice and Ingrid must update their payment channel state to $(a-c,b-d)$. This means Alice deposits $c$ coins and Ingrid deposits $d$ coins for the virtual channel. This update locks the coins ($c+d$) for exclusive use in the virtual channel, while the wardens’ collateral is allocated based on the balance of both the payment channel and the virtual channel. The virtual channel is considered open once the payment channel has been updated accordingly. It’s important to note that opening and closing are the only operations that require changes to the underlying payment channels. Updates to Thunderdome and the payment channels can occur in parallel, as long as the coins locked for the virtual channel are not used during the payment channel update.
@@ 5.2 Thunderdome Update
0:  Parties $A$ and $B$, warden committee $C_{V}$ comprising $C_{A}$ and $C_{B}$, current state $E_{P_{I}}(s_{i})$.
@@
2:  $A$ and $B$ broadcast $M$ to committee $C_{V}$. /*Along with a fee $r$*/
3:  $A$ ($B$) waits for a quorum certificate ($12f+$ 1 signatures) from each warden committee, then $A$ ($B$) execute the virtual channel state $s_{i}$.
@@
0:  Parties $A$ and $B$, warden committee $C_{V}$ comprising $C_{A}$ and $C_{B}$, sequence number $i$.
0:  $C_{V}$ updates to a new valid state.
1:  Each warden $W_{j}$, upon receiving $M$, verifies that both parties’ signatures are present and the sequence number is exactly one higher than the previously stored sequence number. If the warden has published a state or has signed an $M$ or $M^{\prime}$ with the same sequence number, it ignores the state update. Otherwise, $W_{j}$ stores $M$ (as a possible update, not yet replacing the $i-1$-th committed state), and sends its signature $\sigma_{W_{j}}(M)$ to both parties. /*Only to parties that paid the fee*/
@@ 5.3.2 Pessimistic situation
The core idea of Protocol 4 is for the warden committee $C_{V}$ to publish the latest channel state, enabling the closing party to punish malicious wardens by slashing their collateral in the smart contract. At the start of Protocol 4, the closing party submits the register transaction $TX_{r}$ to the smart contract, allowing the payment channel contract to verify published states and proofs-of-fraud using the public keys of the Thunderdome parties and wardens from the respective committees included in $TX_{r}$. With $13f+$ 1 wardens in each committee, the closing party publishes proofs-of-fraud only after at least $12f+$ 1 wardens from their committee have published closing states, ensuring sufficient collateral for penalties.
@@
0:  Party $A$, wardens $W_{1},\cdots,W_{n_{A}}$ from payment channel committee $C_{A}$, virtual channel state $VS$, virtual channel register transaction $TX_{r}$.
@@
1:  $A$ publishes the register transaction $TX_{r}$ on-chain and prove the validity of virtual channel. /*Wardens publish virtual channel states */
@@
3:  After verifying $12f+$ 1 on-chain signed announcements from committee $C_{A}$, $A$ publishes proofs-of-fraud. /*Closing with punishment*/
4:  After the state is included in a block, the smart contract verifies the signature of $VS$ based on the information in the register transaction $TX_{r}$.
@@
6:  Smart contracts award all the cheating wardens’ collateral to party $A$.
7:  First $12f+$ 1 wardens that have no cheating behaviors get an equal fraction of the closing fee.
@@
1:  $SC_{A}$ stores the state with the highest sequence number as $WS_{A}$. If there is more than one, $SC_{A}$ selects the first one it receives. /*Exchange cross-checking transaction*/
2:  $SC_{A}$ sends the notification and query transaction $TX_{qA}$ containning $WS_{A}$ to $SC_{B}$.
3:  $SC_{A}$ waits for the transaction $TX_{qB}$ from $SC_{B}$.
@@
0:  $SC_{A}$ decides on the closing state. /*Receive transaction from the other contract*/
1:  $SC_{A}$ receives transaction $TX_{qB}$ containing $WS_{B}$ from $SC_{B}$. /*React when state is not decided yet*/
2:  If $SC_{A}$ hasn’t deceided $WS_{A}$ yet, then stores $WS_{B}$ as $WS_{A}$ and reply with $NULL$. /*React when state is decided decided*/
@@ 5.4 Multi-hop Thunderdome
The main difference in a multi-hop scenario is the involvement of additional intermediary parties who are unaware of the Thunderdome states. As with the two-hop Thunderdome, it is crucial to ensure that each committee has enough wardens who know the latest Thunderdome state. The key steps for a multi-hop Thunderdome are as follows: (1) During the opening procedure, all four parties, along with the wardens, must collaborate to generate the register transaction. Each party considers the channel open after receiving a quorum certificate from each committee. (2) To successfully update Thunderdome, Alice and Dave must broadcast the updated state to all wardens and wait for a quorum certificate from each committee. (3) Thunderdome can be optimistically closed once all four parties reach an agreement. In a pessimistic closing scenario, each smart contract should wait for at least $12f+$ 1 wardens to publish their states from the corresponding payment channel committee and cross-check with other contracts to determine the final closing state.
@@ 7 Game-theoretic Security Analysis
$Old$ Send a dishonest collaborative closing request with the old virtual channel state
$New$ Send an honest collaborative closing request with the latest virtual channel state
$Uni$ Unilaterally close the part of Thunderdome with Ingrid
@@
$Agree$ Agree and sign the closing request
$Disagree$ Disagree and go for unilateral closing
$Ignore$ Ignore the closing request or go offline
@@ 9 Extensions
Privacy-preserving protocol. In Thunderdome, the privacy of virtual channel states is not protected from the wardens. It is not straightforward to design a privacy-preserving protocol that ensures both: a) Ingrid can recover the state, and b) the wardens cannot retrieve the state, even in collusion with Ingrid, while still being able to verify the validity of the private state. A compromise is to enable state recovery and privacy against wardens, assuming Ingrid does not collude with them. However, even in this case, the repetitive nature of updates poses challenges. Trapdoor one-way functions, though promising candidates, lack IND-CPA security. On the other hand, asymmetric encryption schemes offering such protection require a new random number for each encrypted value [11]. In our protocol, this can be implemented by Alice, Bob, and Ingrid sharing a vector of random values during the creation of the register transaction in Thunderdome open. However, this solution has two main drawbacks: a) the number of states that can be updated in the virtual channel is limited by the length of this vector. Optimistically, Ingrid may occasionally renew it, but this depends on Ingrid being online and cooperative; b) the virtual channel parties must provide the wardens with corresponding evidence (e.g., ZK proof) at each update to certify that the correct random number was used, adding communication and computational overhead to Thunderdome’s update process. Other cryptographic tools, such as Verifiable Secret Sharing [36] or threshold signatures [34], could be leveraged to enhance privacy in Thunderdome transactions. For example, one could explore the possibility of allowing Ingrid and $f+1$ wardens from her committee to recover the states. This scheme may be secure under the assumption of $12f+$ 1 honest wardens, but it remains an open question whether suitable incentive mechanisms can ensure security in the rational model.

### v4

@@ 1.2 Our contribution
  • We formalize the security properties of Thunderdome and conduct a security analysis within the honest/Byzantine model (Section 6). Specifically, we demonstrate that our protocol is secure under the assumption of distrusting participants, where $f$ out of $3f+1$ wardens in each payment channel are Byzantine, and the remainder are honest.
@@ 2.2 Brick channel
Channel structure. In addition to the two primary parties, Alice and Bob, a Brick channel includes a committee of $3f+1$ wardens, where up to $f$ wardens can behave maliciously (Byzantine). The basic structure of the Brick channel is illustrated in Fig. 1.
@@
Brick open. Once Alice and Bob agree to open a Brick channel, they broadcast the channel information to the wardens. Each party can consider the channel open only after receiving at least $2f+1$ signed acknowledgments from the wardens – a quorum certificate. The threshold of $2f+1$ ensures safety in asynchronous communication networks.
Brick update. Each state generated by Alice and Bob is hashed and signed by both parties. The state is then assigned a sequence number to indicate its order. This sequence number, along with the hashed state, is signed and broadcast to the wardens. Once at least $2f+1$ wardens have signed the sequence number, the state is considered valid, and the parties can execute it.
Brick close. The Brick channel closing protocol has two scenarios: optimistic and pessimistic. In the optimistic case, both parties are online and responsive, allowing them to collaboratively generate and sign a closing request, which is then published on-chain. In the pessimistic case, one party, say Alice, may be offline for an extended period and unresponsive to Bob’s close request. To prevent the channel from being indefinitely locked, the Brick protocol allows Bob to close the channel unilaterally. However, Bob alone does not have a valid closing request, as during the channel update, the parties only exchange signatures on the hashed state, which is insufficient to close the channel. To close unilaterally, Bob initiates a closing request on-chain, prompting the wardens to publish the latest stored sequence number. The valid closing state is defined by the highest sequence number signed by Alice, Bob, and at least one warden. To incentivize honest behavior from the wardens, Brick employs a punishment mechanism based on proofs-of-fraud. A proof-of-fraud consists of a warden’s signature on an update with a higher sequence number than the one they submitted for the closing request. Bob collects these signatures during the channel update process. If a valid proof-of-fraud is provided, Bob can claim the corresponding warden’s collateral on-chain. According to Brick’s security analysis, if the aggregate collateral of wardens submitting outdated states exceeds the channel balance, Bob has no incentive to close the channel incorrectly. Instead, he would claim the wardens’ collateral and award the entire channel balance to his counterparty. Specifically, for a Brick channel with $v$ coins as the balance, each of the $3f+1$ wardens must deposit at least $\frac{v}{f}$ coins on-chain to ensure security.
@@ 3.2 Threat models
Byzantine security model. In this model, we assume each committee consists of $3f+1$ wardens, of which $f$ may be Byzantine, while the rest are honest and follow the protocol as specified. We note that the protocol’s security can be extended to scenarios where the number of wardens differs as long as the Byzantine ratio remains no larger than $\frac{f}{3f+1}$ (i.e., the total number of wardens may vary depending on the setup, but the Byzantine proportion must be maintained). In addition to wardens, the security model also accounts for cases where the main Thunderdome parties may be Byzantine, ensuring security for any honest participant. This is the weakest meaningful assumption, as collusion between all parties against a single honest party could occur; however, considering all parties deviating from the protocol holds no value since security properties apply only to honest participants. Byzantine nodes can arbitrarily deviate by delaying, crashing, or signing/publishing incorrect states or messages. However, they cannot drop honest messages or forge signatures, as these actions violate network and cryptographic assumptions.
@@ 4 Protocol Overview
Naive design. In a naive design of Thunderdome, we follow the same rationale as regular TPC protocols like Brick. When opening the virtual channel, the wardens of each TPC lock a portion of their collateral into Thunderdome. For example, if 2 out of 10 coins in a TPC are locked in the virtual channel, each warden locks 20% of their collateral during the funding transaction. The update phase of Thunderdome operates similarly to Brick. However, the closing phase differs during collaborative closing: Thunderdome cannot be closed solely by Alice and Bob; Ingrid’s agreement is also required. For unilateral closing, the smart contract only requires a quorum certificate ($2f+1$ signatures) from the closing party’s warden committee (e.g., $C_{A}$ or $C_{B}$).
@@
Once we establish how smart contracts should handle states published by wardens, the next question is how many on-chain publications are necessary, given that asynchrony only guarantees eventual message delivery. If an updated state is confirmed with only $2f+1$ signatures from each warden committee, there could still be up to $f$ (honest) wardens who have not received the latest state, even after the protocol was closed (incorrectly). To mitigate this, we require the smart contract to wait for at least $f+1$ publishments from the corresponding committee before closing the channel. This ensures that at least one warden with locked collateral knows the latest state.
@@
Thus, we require at least $2f+1$ wardens from each committee, $C_{A}$ and $C_{B}$, to publish their closing state in Thunderdome’s unilateral closing protocol. Combined with the quorum certificates required to update the Thunderdome state, we ensure that at most $f$ wardens per committee can be slow. By waiting for $2f+1$ publications from accountable wardens, at least $f+1$ publications can be punished if they submit outdated information. Given that each warden locks $v/f$ coins in Thunderdome, the total collateral that can be claimed per channel in case of fraud is $(f+1)\times\frac{v}{f}>v$.
@@ 5.1 Thunderdome Open
0: Parties AA, II and $B$, TPC committees $C_{A}$ and $C_{B}$, initial virtual channel state $s_{1}$, total channel balance $v$ and contract information cici.
@@
1:AA, II and $B$ combine wardens from $C_{A}$ and $C_{B}$ to form virtual channel committee $C_{V}=C_{A}\cup C_{B}$. /*Parties reach agreement on the open state*/
2:AA and $B$ generate two pre-register transaction with their signatures $TX_{r}^{\prime}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A}/\sigma_{B}\}$ and send to $I$.
3: After receiving two transactions, II responds with $TX_{r}^{\prime\prime}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A}/\sigma_{B},\sigma_{I}\}$.
4:AA and $B$ exchange responses of $I$ and generate final register transaction with all partis’ signatures $TX_{r}=\{pk_{A},pk_{B},pk_{I},\{pk_{C_{V}}\},s_{1},v,ci,\sigma_{A},\sigma_{B},\sigma_{I}\}$.
@@
6:AA and II (II and $B$) update payment channels according to initial virtual channel state $s_{1}$.
@@
The final step in opening Thunderdome, after all parties reach an agreement, is to "virtually lock" the coins for the virtual channel. Unlike payment channels, which require publishing funding transactions on-chain, the parties of both underlying payment channels—Alice and Ingrid, and Ingrid and Bob—update their respective payment channels to lock coins for use in the virtual channel. For example, if the payment channel state between Alice and Ingrid is $(a,b)$, and Alice wants to open a virtual channel with Bob with an initial state of $(c,d)$, then Alice and Ingrid must update their payment channel state to (ac,bd)(a-c,b-d). This means Alice deposits $c$ coins and Ingrid deposits $d$ coins for the virtual channel. This update locks the coins ($c+d$) for exclusive use in the virtual channel, while the wardens’ collateral is allocated based on the balance of both the payment channel and the virtual channel. The virtual channel is considered open once the payment channel has been updated accordingly. It’s important to note that opening and closing are the only operations that require changes to the underlying payment channels. Updates to Thunderdome and the payment channels can occur in parallel, as long as the coins locked for the virtual channel are not used during the payment channel update.
@@ 5.2 Thunderdome Update
0: Parties AA and $B$, warden committee $C_{V}$ comprising $C_{A}$ and $C_{B}$, current state $E_{P_{I}}(s_{i})$.
@@
2:AA and $B$ broadcast $M$ to committee $C_{V}$. /*Along with a fee rr*/
3:AA ($B$) waits for a quorum certificate ($2f+1$ signatures) from each warden committee, then AA ($B$) execute the virtual channel state $s_{i}$.
@@
0: Parties AA and $B$, warden committee $C_{V}$ comprising $C_{A}$ and $C_{B}$, sequence number ii.
[…]
1: Each warden $W_{j}$, upon receiving $M$, verifies that both parties’ signatures are present and the sequence number is exactly one higher than the previously stored sequence number. If the warden has published a state or has signed an $M$ or $M^{\prime}$ with the same sequence number, it ignores the state update. Otherwise, $W_{j}$ stores $M$ (as a possible update, not yet replacing the i1i-1-th committed state), and sends its signature $\sigma_{W_{j}}(M)$ to both parties. /*Only to parties that paid the fee*/
@@ 5.3.2 Pessimistic situation
The core idea of Protocol 4 is for the warden committee $C_{V}$ to publish the latest channel state, enabling the closing party to punish malicious wardens by slashing their collateral in the smart contract. At the start of Protocol 4, the closing party submits the register transaction $TX_{r}$ to the smart contract, allowing the payment channel contract to verify published states and proofs-of-fraud using the public keys of the Thunderdome parties and wardens from the respective committees included in $TX_{r}$. With $3f+1$ wardens in each committee, the closing party publishes proofs-of-fraud only after at least $2f+1$ wardens from their committee have published closing states, ensuring sufficient collateral for penalties.
@@
0: Party AA, wardens $W_{1},\cdots,W_{n_{A}}$ from payment channel committee $C_{A}$, virtual channel state $VS$, virtual channel register transaction $TX_{r}$.
@@
1:AA publishes the register transaction $TX_{r}$ on-chain and prove the validity of virtual channel. /*Wardens publish virtual channel states */
@@
3: After verifying $2f+1$ on-chain signed announcements from committee $C_{A}$, AA publishes proofs-of-fraud. /*Closing with punishment*/
4: After the state is included in a block, the smart contract verifies the signature of VSVS based on the information in the register transaction $TX_{r}$.
@@
6: Smart contracts award all the cheating wardens’ collateral to party AA.
7: First $2f+1$ wardens that have no cheating behaviors get an equal fraction of the closing fee.
@@
[…]
@@
[…]
2: If $SC_{A}$ hasn’t deceided $WS_{A}$ yet, then stores $WS_{B}$ as $WS_{A}$ and reply with NULLNULL. /*React when state is decided decided*/
@@ 5.4 Multi-hop Thunderdome
The main difference in a multi-hop scenario is the involvement of additional intermediary parties who are unaware of the Thunderdome states. As with the two-hop Thunderdome, it is crucial to ensure that each committee has enough wardens who know the latest Thunderdome state. The key steps for a multi-hop Thunderdome are as follows: (1) During the opening procedure, all four parties, along with the wardens, must collaborate to generate the register transaction. Each party considers the channel open after receiving a quorum certificate from each committee. (2) To successfully update Thunderdome, Alice and Dave must broadcast the updated state to all wardens and wait for a quorum certificate from each committee. (3) Thunderdome can be optimistically closed once all four parties reach an agreement. In a pessimistic closing scenario, each smart contract should wait for at least $2f+1$ wardens to publish their states from the corresponding payment channel committee and cross-check with other contracts to determine the final closing state.
@@ 7 Game-theoretic Security Analysis
OldOld Send a dishonest collaborative closing request with the old virtual channel state
NewNew Send an honest collaborative closing request with the latest virtual channel state
UniUni Unilaterally close the part of Thunderdome with Ingrid
@@
AgreeAgree Agree and sign the closing request
DisagreeDisagree Disagree and go for unilateral closing
IgnoreIgnore Ignore the closing request or go offline
@@ 9 Extensions
Privacy-preserving protocol. In Thunderdome, the privacy of virtual channel states is not protected from the wardens. It is not straightforward to design a privacy-preserving protocol that ensures both: a) Ingrid can recover the state, and b) the wardens cannot retrieve the state, even in collusion with Ingrid, while still being able to verify the validity of the private state. A compromise is to enable state recovery and privacy against wardens, assuming Ingrid does not collude with them. However, even in this case, the repetitive nature of updates poses challenges. Trapdoor one-way functions, though promising candidates, lack IND-CPA security. On the other hand, asymmetric encryption schemes offering such protection require a new random number for each encrypted value [11]. In our protocol, this can be implemented by Alice, Bob, and Ingrid sharing a vector of random values during the creation of the register transaction in Thunderdome open. However, this solution has two main drawbacks: a) the number of states that can be updated in the virtual channel is limited by the length of this vector. Optimistically, Ingrid may occasionally renew it, but this depends on Ingrid being online and cooperative; b) the virtual channel parties must provide the wardens with corresponding evidence (e.g., ZK proof) at each update to certify that the correct random number was used, adding communication and computational overhead to Thunderdome’s update process. Other cryptographic tools, such as Verifiable Secret Sharing [36] or threshold signatures [34], could be leveraged to enhance privacy in Thunderdome transactions. For example, one could explore the possibility of allowing Ingrid and $f+1$ wardens from her committee to recover the states. This scheme may be secure under the assumption of $2f+1$ honest wardens, but it remains an open question whether suitable incentive mechanisms can ensure security in the rational model.
@@
Acknowledgement
This work was supported by the Vienna Science and Technology Fund (WWTF) through the project 10.47379/ICT22045;